# Great Models Think Alike and this Undermines AI Oversight

### v1

@@ 7 Conclusion, Limitations, Future Work
Our paper shows the importance of measuring functional similarity for language models. We derive a novel, probabilistic metric for model similarity, CAPA ($\kappa_{p}$). We then use it to study the implications of similarity for AI oversight – showing affinity bias in AI judges, and the role of complementary knowledge when training on LM annotations, such as in weak-to-strong generalization. AI oversight will become more relevant as capabilities improve, so our finding that increasing capabilities could lead to more correlated errors is particularly concerning. Thus, we believe measuring and accounting for model similarity is going to be increasingly important. We now list some limitations of our work, along with avenues for future work that can help develop a better understanding of model similarity and its implications.
@@
Extending to similarity metrics for free-text outputs: Everyday use of generative models is based on their free-text responses. Like much work on benchmarking, we had to limit to MCQ tasks as the science of precisely evaluating free-text is still evolving (Biderman et al., 2024). For example, both model-free (Papineni et al., 2002) and model-based metrics (Pillutla et al., 2021) suffer from a wide range of syntax and style sensitivity issues (Kocmi et al., 2021; He et al., 2023). We hope the community takes up the challenge of designing similarity metrics for free-response text and reasoning. This would allow studying the role of similarity using more promising oversight setups like debate (Kenton et al., 2024) and process supervision (Lightman et al., 2023).
Generator-Verifier gap: AI oversight has recently shown promise for tasks where it is easier to validate a solution than generate it (Song et al., 2024). Similarity may continue to play a role here. (1) In evaluations, similarity in stylistic preferences of the generated solution may influence judge scores. (2) In training, the generator-verifier gap may be larger if models are more different.
Safety implications: Researchers separately develop many post-training interventions to reduce harmfulness, dual-use knowledge, dishonesty etc. In real world model deployments, all these problems have to be tackled at once, which can benefit from composing interventions (Kolbeinsson et al., 2024). If the benefits are decorrelated, composing would lead to greater compound safety. If the side effects are correlated, composing would lead to lower accumulation. More broadly, measuring LM similarity post-intervention can help characterize decorrelation in research bets, for granters (Canton, 2025). For example, LM unlearning was recently found to be functionally similar to refusal training (Łucki et al., 2024), even though it was proposed as a complementary safeguard (Li et al., 2024a). Finally, as we transition towards language agents, similarity can help understand collective “blind spots” (He et al., 2023), and could lead to better cooperation (Lowe et al., 2017) but also scheming (Balesni et al., 2024) between multiple agents.
Qualitative analysis of model differences: We developed quantitative methods for measuring LM similarity on a given data distribution. One exciting direction is to use these metrics to provide qualitative difference descriptors (Dunlap et al., 2024a) between models, by describing data distributions where models are least similar.
@@ A.2 Extending CAPA to more than two models
A.3 How to use CAPA for classification and exact match settings?
@@ A.4 Detailed Discussion on Design Choices
Why not use Matthews Correlation Coefficient: We could take a completely different approach by computing the Pearson or Matthews Correlation Coefficient of the binary vectors of sample-wise correctness for the two models (Chicco et al., ). However, it would be difficult to incorporate probabilistic information, and that models can be incorrect and still disagree by predicting different options. In other words, it suffers from the same issues as error consistency, and we found it more difficult to extend.
@@ A.7 CAPA comparison with other inter-rater metrics
Furthermore, we design a simulation experiment to compare the ’behavior’ of the above listed inter-rater metrics with our novel contribution CAPA. In particular, we limit the simulation to a binary classification problem as standard metrics like Cohen’s $k$ and Scott’s $\pi$ are ill-suited for multiple choice question settings. In total we investigate the performance of 4 metrics: Cohen’s $k$ Probabilistic, Scott’s $\pi$ Probabilistic, Error consistency, and CAPA. We simulate N=10000 observations for 2 models. For the first model we set it’s accuracy to 90%, it always favors the 1st option, and has a high calibration, 0.99, meaning the model is highly confident in it’s predictions (e.g. single data point is [0.99, 0.01]). For the second model we iteratively increase it’s accuracy by adjusting it’s calibration from 0.01 to 0.99 for the first option, as such, making the models more similar artificially.
The first observation from the results reported in Fig. 6 is that both standard inter-rater metrics, Cohen’s $\kappa$ and Scott’s $\pi$ (even when adjusted to take into account probabilities) are ill suited for the present use-case: capturing model similarity. The main issue stems from the fact that the computation of $P_{e}$ if simply adjusted to probabilistic setting without taking into account model accuracy, obtains a similar computational value as $P_{o}$ (in this case equal to $c_{\text{obs}}^{p}$). $P_{e}^{p}$ computes marginal class probabilities as indicated in Section A.5, which is ill suited when the model attributes all it’s probability mass to a single option (always prefers option A in MCQ setting). Furthermore, whilst error consistency improves upon Cohen’s $\kappa$ and Scott’s $\pi$ it over estimates model similarity. In particular, when both models reach 90% accuracy error consistency reports perfect agreement, while in reality model output probabilities differ, [0.99, 0.01] and [0.65, 0.35] respectively. As such, our metric is the only one that is able to capture model observed agreement $c_{\text{obs}}^{p}$ increasing beyond model accuracy levels and reaching 1 when models are highly calibrated, e.g. [0.99, 0.01] and [0.99, 0.01].
@@
Figure 6: Metric comparison when models tend towards agreement. We compare different metric values for two models in a binary setting. For first model we set 90% accuracy and calibration to 0.99 (meaning the model is highly confident in its answers). For the second model, we increase it’s calibration from 0.01 to 0.99 to approach the same distribution as the first model. On y-axis we are plotting metric value on x-axis we are reporting $\overline{p}$ for the second model which as the model becomes more calibrated approaches accuracy of the first model.
Figure 7: Metric comparison when models tend towards disagreement (Read plot from right to left). We compare different metric values for two models in a binary setting. For the first model, we set accuracy to 90% and calibration to 0.99 (the model is highly confident in its answers). For the second model, we incrementally increase its disagreement with model one by pushing its probability mass to the second option and increasing its calibration to 0.99.
Figure 8: Metric comparison when models tend towards disagreement with adjusted $\kappa_{p}$. Replication of fig. 8 but with adjusted $\kappa_{p}$ as as $\hat{\kappa_{p}}$, computation following eq. 18.
@@
In addition, we investigated the ’behavior’ of the above listed metrics as the models become increasingly dissimilar. In this set up we change that the second model always prefers the second option. Thus, by iteratively increasing its calibration we obtain models that maximally differ in their probability distribution, e.g. [0.99, 0.01] and [0.01, 0.99] respectively. As such, also the accuracy of the second model decreases overtime from random chance (50%) to 0.10 %, and we would like to obtain metric of -1. As it can be seen in Fig. 8, CAPA never reaches -1. Importantly, the same issue also can be observed for error consistency. This observation comes from the fact that both metrics use the original Cohen’s $\kappa$ equation. As explained in Section A.6, $\kappa_{p}=-1$ iff $c_{\text{exp}}^{p}=0.5$. For probabilistic Cohen’s $\kappa$ we see the same observation as in Fig. 6, the marginal probability computation is not suited for the given problem. Interestingly, probabilistic Scott’s $\pi$ is the only metric that approaches -1. Whilst a desired final outcome, Scott’s $\pi$ overestimates model disagreement when model probabilities are independent, [0.99, 0.01] and [0.5, 0.5].
@@ Appendix B LLM-as-a-Judge
B.1 Comparison of Judge Scores for Our Similarity vs Error Consistency
@@
B.2 Evaluating Judge Scores Against Ground-Truth
B.2.1 MCQ Ground-Truth
@@
B.2.2 Judge Ensemble with Access to Reference Answers
@@
Experimental Setup Every response is evaluated using an ensemble of five capable LMs used as LLM-as-a-judge from a range of different model families. The judge is given access to the question, the model’s free-form response and all MMLU-Pro reference options. For each option we indicate if it is the correct or a wrong option. Using this information, the judge has to decide whether the model’s response is correct or wrong. The prompt can be seen in Prompt B.8.2. For every per-sample response, we aggregate the five binary decisions using majority voting. Since there are five judges and it is a binary decision task, there are no ties. A qualitative analysis has shown the high quality of this process in determining the correctness of responses. The judges used are gemma-2-27b-it, Qwen2.5-32B-Instruct, Qwen2.5-72B-Instruct, Llama-3.1-70-Instruct and Llama-3.3-70B-Instruct (Gemma Team, 2024; Qwen Team, 2025; Llama Team, 2024a, c).
@@
B.2.3 Judge Score Validity Against Reference-Based Ensemble
@@
Capability-Dependent Affinity Effects Even if the ordinal ranking of evaluated models is mostly accurate, there is a consistent trend that too many wrong responses are judged as being right. The exact behavior varies from judge to judge. Consider the small Llama-3.1-8B-Instruct for instance: it has a consistent positivity bias and ranks too many wrong responses as correct, even for models of low capability. Qwen2.5-72B-Instruct on the other hand appears to be much more capable in identifying the wrong responses of low-capability models. However, as the evaluated LMs become stronger, it exhibits the same bias as the smaller Llama judge. This aligns with the findings of Section 3 that LLM-as-judges show an affinity bias, because more capable models are also more similar to Qwen2.5-72B-Instruct.
@@
B.3 Statistical Testing

### v2

@@ 7 Conclusion, Limitations, Future Work
Our paper shows the importance of measuring functional similarity for language models. We derive a novel, probabilistic metric for model similarity, CAPA ($\kappa_{p}$). We then use it to study the implications of similarity for AI oversight – showing affinity bias in AI judges, and the role of complementary knowledge when training on LM annotations, such as in weak-to-strong generalization. AI oversight will become more relevant as capabilities improve, so our finding that increasing capabilities could lead to more correlated errors is particularly concerning. Thus, we believe measuring and accounting for model similarity is going to be increasingly important. We now list some limitations of our work and avenues for future work, that can help develop a better understanding of model similarity and its implications.
@@
Extending to similarity metrics for free-text outputs: Everyday use of generative models is based on their free-text responses. Like much work on benchmarking, we had to limit to MCQ tasks as the science of precisely evaluating free-text is still evolving (Biderman et al., 2024). For example, both model-free (Papineni et al., 2002) and model-based metrics (Pillutla et al., 2021) suffer from a wide range of syntax and style sensitivity issues (Kocmi et al., 2021; He et al., 2023). We hope the community takes up the challenge of designing similarity metrics for free-response text and reasoning. This would allow studying the role of similarity using more promising oversight setups like debate (Kenton et al., 2024) and process supervision (Lightman et al., 2024).
Generator-Verifier gap: AI oversight has recently shown promise for tasks where it is easier to validate a solution than generate it (Song et al., 2025). Similarity may continue to play a role here. (1) In evaluations, similarity in stylistic preferences of the generated solution may influence judge scores. (2) In training, the generator-verifier gap may be larger if models are more different.
Safety implications: Researchers separately develop many post-training interventions to reduce harmfulness, dual-use knowledge, dishonesty etc. In real world model deployments, all these problems have to be tackled at once, which can benefit from composing interventions (Kolbeinsson et al., 2025). If the benefits are decorrelated, composing would lead to greater compound safety. If the side effects are correlated, composing would lead to lower accumulation. More broadly, measuring LM similarity post-intervention can help characterize decorrelation in research bets, for granters (Canton, 2025). For example, LM unlearning was recently found to be functionally similar to refusal training (Łucki et al., 2025), even though it was proposed as a complementary safeguard (Li et al., 2024a). Finally, as we transition towards language agents, similarity can help understand collective “blind spots” (He et al., 2023), and could lead to better cooperation (Lowe et al., 2017) but also scheming (Balesni et al., 2024) between multiple agents.
Qualitative analysis of model differences: We developed quantitative methods for measuring LM similarity on a given data distribution. One exciting direction is to use these metrics to provide qualitative difference descriptors (Dunlap et al., 2025) between models, by describing data distributions where models are least similar.
@@ Acknowledgments
Impact Statement
Our paper can be considered as foundational research analyzing mechanisms of AI oversight and identifying an important shortcoming, and not tied to particular applications.
@@ A.2 Extending CAPA to more than two models
A.3 How to use CAPA beyond multiple choice?
@@
Regression: For tasks with numeric predictions like regression, for the observed agreement $c_{\text{obs}}$, one could measure a distance metric over the two models’ predictions, aggregating across samples. Once again, models with lower error would have lower distance in prediction, so the challenge lies in defining chance agreement $c_{\text{exp}}$, for a model with a given error. We would have to make appropriate assumptions about the distribution of errors, such as assuming they are gaussian, and compute $c_{\text{exp}}$ accordingly.
@@ A.4 Detailed Discussion on Design Choices
Why not use Matthews Correlation Coefficient: We could take a completely different approach by computing the Pearson or Matthews Correlation Coefficient of the binary vectors of sample-wise correctness for the two models (Chicco et al., 2021). However, it would be difficult to incorporate probabilistic information, and that models can be incorrect and still disagree by predicting different options. In other words, it suffers from the same issues as error consistency, and we found it more difficult to extend.
@@
Why a uniform prior? In the present work we make the assumption that the prior error distribution follows a uniform distribution. In the context of multiple choice tasks, where by design each answer should have an equal distribution in the dataset, this is a valid and data grounded assumption. However, our metric is not limited to a uniform prior assumption. In principle, one can easily adjust the agreement by chance calculation (eq. 10) by a different prior distribution assumption given the data observed, for example, a categorical distribution if the class distribution in the data is non-uniform.
@@ A.7 CAPA comparison with other inter-rater metrics
Simulation Experiments We design a simulation set-up to compare the ’behavior’ of the above listed inter-rater metrics with our novel contribution CAPA. In particular, we limit the simulation to a binary classification problem as standard metrics like Cohen’s $k$ and Scott’s $\pi$ are ill-suited for multiple choice question settings. In total we investigate the performance of 5 metrics: Cohen’s $k$ Probabilistic, Scott’s $\pi$ Probabilistic, Error consistency, JS-Distance and CAPA. We simulate N=10000 observations for 2 models.
First, we investigate the metric behavior when both models tend towards agreement and their errors are correlated. In particular, we set the accuracy of model A to 90% and it always favors the 1st option (has a high calibration, 0.99), meaning the model is highly confident in it’s predictions (e.g. single data point is [0.99, 0.01]). For the other model, model B, we iteratively increase it’s accuracy by adjusting it’s calibration from 0.01 to 0.99 for the first option, as such, making the models more similar artificially.
@@
Figure 7: Metric comparison when two models tend towards agreement. For model A we set accuracy to 90% and calibration to 0.99 of the correct answer. For model B we increase its calibration from 0.01 to 0.99 of the correct answer, so its answer distribution becomes increasingly similar to model A. On y-axis we are plotting metric value on x-axis we are reporting $\overline{p}$ for model B which, as the model becomes more calibrated, approaches accuracy of model A.
As can be noted in Fig. 7 all previous similarity metrics fail to capture the full range of increasing model similarity: (1) EC only captures model similarity up to when their accuracies are matched (both models reach 90% accuracy), yet fails to further distinguish between models that have high agreement [0.99, 0.01] and [0.99, 0.01] versus medium [0.99, 0.01] and [0.75, 0.25]; (2) while Jensen–Shannon distance (computed as 1 - JS-distance) correctly captures the gradual increase in model similarity it overestimates it initially, this is due to the fact that the distance is computed from the average joint distribution; (3) the probabilistic counterparts of standard inter-rater metrics like Cohen’s $\kappa$ and Scott’s $\pi$ (derivation details in App. A.5) are also ill suited. In particular, in the case of Cohen’s $\kappa$ prob the class marginals (for computing the expected agreement) match the observed agreement, therefore leading to inter-rater agreement of 0 even with increasing model similarity, while in the case of Scott’s $\pi$ prob we do observe a change in similarity score due to it’s assumption of joint marginal distribution, but also because of this assumption the metric overestimates the expected agreement leading to negative (opposing) inter-rater agreement, which is an incorrect conclusion given the simulation data.
Second, we explore the metric behavior when both model accuracies are increasing yet errors remain random (uncorrelated). We again simulate 10’000 binary data points with uniform class distribution. For model A we set the accuracy to 90%, while for model B we incrementally (100 steps) increase it’s accuracy from 50% to 90% by randomly selecting data samples which are correct at each step. As such, the errors are uncorrelated between the two models. Results are reported in main paper in Sec. A.7.
@@
In addition, we perform a simulation experiment when the two models become increasingly dissimilar. In this set up we change that the model B always prefers the second option. Thus, by iteratively increasing its calibration we obtain models that maximally differ in their probability distribution, e.g. [0.99, 0.01] and [0.01, 0.99] respectively. As such, also the accuracy of model B decreases overtime from random chance (50%) to 0.10 %, and we would like to obtain metric of -1. As it can be seen in Fig. 9, CAPA never reaches -1. Importantly, the same issue also can be observed for error consistency. This observation comes from the fact that both metrics use the original Cohen’s $\kappa$ equation. As explained in Section A.6, $\kappa_{p}=-1$ iff $c_{\text{exp}}^{p}=0.5$. For probabilistic Cohen’s $\kappa$ we see the same observation as in Fig. 7, the marginal probability computation is not suited for the given problem. Interestingly, probabilistic Scott’s $\pi$ is the only metric that approaches -1. Whilst a desired final outcome, Scott’s $\pi$ overestimates model disagreement when model probabilities are independent, [0.99, 0.01] and [0.5, 0.5].
@@
Figure 8: Metric comparison when models tend towards disagreement (Read plot from right to left). We compare different metric values for two models in a binary setting. For model A, we set accuracy to 90% and calibration to 0.99 (the model is highly confident in its answers). For model B, we incrementally increase its disagreement with model A by pushing its probability mass to the second option and increasing its calibration to 0.99.
Figure 9: Metric comparison when models tend towards disagreement with adjusted CAPA. Replication of fig. 9 but with adjusted CAPA as $\hat{\kappa_{p}}$, computation following eq. 18.
@@ Appendix B LLM-as-a-Judge
B.1 Extended Multiple Regression Analysis with Model Size as a Confounder
We extend the main paper regression analysis (Sec. 3) by also including model size as a possible confounder of judgment score. We standardize model size by removing the mean and scaling to unit variance, so the model size as a independent variable is on a comparable scale as model similarity and accuracy. We remove 6 outliers based on the fact that the studentized residuals $>3$. All outliers where in combination with model ’HuggingFaceTB-SmolLM2-135M-Instruct’. The extended regression results are reported in table 6. We report that even when controlling for model accuracy and model size, our computed model similarity score remains a significant predictor of judgment score. Furthermore, whilst for some models model size is a significant predictor of judgment score, when controlling for model similarity and accuracy, the effect size is very small, around 0. All in all, the additional results confirm that our metric accounts for model size as a possible confounder.
Table 6: Extended Multiple Regression Results
Judge sim acc size
Qwen2.5-7B-It 0.49∗∗ 0.68∗∗ -0.02
Qwen2.5-32B-It 0.40∗∗ 1.02∗∗ -0.03∗∗
Qwen2.5-72B-It 0.42∗∗ 1.23∗∗ -0.03∗∗
Meta-Llama-3.1-8B-It 0.93∗∗ 0.61∗∗ -0.02
Meta-Llama-3.1-70B-It 0.74∗∗ 0.92∗∗ -0.03∗∗
Llama-3.3-70B-It 0.61∗∗ 1.03∗∗ -0.04∗∗
gemma-2-9b-It 0.67∗∗ 0.78∗∗ -0.02
gemma-2-27b-It 0.62∗∗ 0.77∗∗ -0.02
Ministral-8B-It-2410 0.65∗∗ 0.50∗∗ -0.02
B.2 Comparison of Judge Scores for Our Similarity vs Error Consistency
@@
B.3 Evaluating Judge Scores Against Ground-Truth
B.3.1 MCQ Ground-Truth
@@
B.3.2 Judge Ensemble with Access to Reference Answers
@@
Experimental Setup Every response is evaluated using an ensemble of five capable LMs used as LLM-as-a-judge from a range of different model families. The judge is given access to the question, the model’s free-form response and all MMLU-Pro reference options. For each option we indicate if it is the correct or a wrong option. Using this information, the judge has to decide whether the model’s response is correct or wrong. The prompt can be seen in Prompt B.11.2. For every per-sample response, we aggregate the five binary decisions using majority voting. Since there are five judges and it is a binary decision task, there are no ties. A qualitative analysis has shown the high quality of this process in determining the correctness of responses. The judges used are gemma-2-27b-it, Qwen2.5-32B-Instruct, Qwen2.5-72B-Instruct, Llama-3.1-70-Instruct and Llama-3.3-70B-Instruct (Gemma Team, 2024; Qwen Team, 2025; Llama Team, 2024a, c).
@@
B.3.3 Judge Score Validity Against Reference-Based Ensemble
@@
Capability-Dependent Affinity Effects Even if the ordinal ranking of evaluated models is mostly accurate, there is a consistent trend that too many wrong responses are judged as being right. The exact behavior varies from judge to judge. Consider the small Llama-3.1-8B-Instruct for instance: it has a consistent positivity bias and rates too many wrong responses as correct, even for models of low capability. Qwen2.5-72B-Instruct on the other hand appears to be much more capable in identifying the wrong responses of low-capability models. However, as the evaluated LMs become stronger, it exhibits the same bias as the smaller Llama judge. This aligns with the findings of Section 3 that LLM-as-judges show an affinity bias, because more capable models are also more similar to Qwen2.5-72B-Instruct.
@@
B.4 Judge Preference in Chat Generation Tasks Against Similarity
Experimental Setup To evaluate judge bias in a different context, we evaluated each of the 23 instruction-tuned models of the original experiment on the benchmark AlpacaEval for free-form generation (Dubois et al., 2023). After generating responses for all models on the full set of AlpacaFarm prompts, we used the same nine LMs as judges as before to determine their preferences on all response pairs between the 253 model combinations. A re-implementation of AlpacaEval 2.0 was used to determine preferences (Dubois et al., 2024). The resulting judge preferences were shuffled and used to compute Judgment Elo for each judge. In total, there were 203,665 combinations of prompts and pairs of models to be judged.
Free-form Judge Preference Figure 13 shows the relation between a model’s Judgment Elo on Alpaca Eval and judge-model similarities using our proposed metric CAPA on MMLU-Pro. Since Alpaca Eval cannot be easily turned into an MCQ benchmark, as required by our similarity metric, using the same similarities on MMLU-Pro as before was inevitable. Despite the domain gap between the data used for the LLM-as-a-judge experiments and CAPA, there is a clear positive correlation between the two, indicating the existence of an affinity bias. This also provides initial evidence that similarity on a diverse MCQ benchmark might transfer as a predictor across tasks.
Figure 13: Elo of LMs on AlpacaEval 2.0 as judged by different LLM-as-judges in relation with our similarity metric on MMLU-Pro. We compare the Elo obtained by using different judges for binary preference pairs on Alpaca Eval to their similarity on the filtered MMLU-Pro dataset.
B.5 Judge Score Against Similarity per Category
MMLU-Pro is a very comprehensive MCQ benchmark to evaluate a model’s common knowledge. By filtering some of the questions to ensure that only those that can be answered in free-form remain, we may have unintentionally introduced biases such as Simpson’s paradox that affect the preference bias we observed. To ensure that this is not the case, we group the questions for free-form Judgment Scores and MCQ similarity by their original category in Figure 14.
For each category, we only used the judgments and MCQ responses of the questions belonging to that category to compute the Judgment Scores and CAPA as metric for model similarity. As in Figure 3, each line corresponds to one judge model. The same positive correlation between CAPA and Judgment Score can be seen for every category as on the original, aggregated Filtered MMLU-Pro benchmark. This indicates the existence of an affinity bias of LLM-as-a-judge irrespective of the category.
For better readability, we do not report the Pearson r in individual legends for each subplot. Apart from one exception, all Pearson correlation coefficients of the LLM-as-judges in each category were above 0.75 and statistically significant (p=0.01). The only exception was google/gemma-2-9b-it on the category “engineering” with $r=0.28$ and $p=0.211$.
Figure 14: Judgment Score Relation with Model Similarity per Category. This plot shows the same experiment as Figure 3 except that all questions were grouped by MMLU-Pro category. Each line is a regression model fit between judgment and similarity scores as computed between model and judge pairs. Each point represents a single pair, and $\diamond$ indicates that both, the model and the judge, come from the same model family.
B.6 Statistical Testing